OpenStudy (anonymous):

the sum of the squares of two consecutive positive odd inteegers is 202 find the numbers

OpenStudy (anonymous):

guess and check

OpenStudy (anonymous):

i did and it didnt add up

OpenStudy (anonymous):

\(7^2+9^2=49+81\) is too small, try \(9^2+11^2\)

OpenStudy (anonymous):

also let x = number 1 let number 2 = x +2 x^2+(x+2)^2=202 the positive root will give you x=9, therefore the numbers are 9 and 11

OpenStudy (anonymous):

x^2 + (x +2)^2 = 202